# Complexity of Abduction in Łukasiewicz Logic

### v1

@@ Appendix C Proofs of Section 5
Now let $\mathbf{v}$ be a classical valuation s.t. $\mathbf{v}(\phi)=1$ for all $\phi\in\Gamma_{\mathcal{H}}$, $\mathbf{v}(\sigma_{\mathcal{H}})=1$ but $\mathbf{v}(\tau_{\mathcal{H}})=0$. We need to define an Ł-valuation $v_{\mathbf{v}}$ s.t. $v_{\mathbf{v}}(\lambda)=1$ iff $\mathbf{v}(r_{\lambda})=1$ would hold for every $\lambda$ occurring in $\mathbb{P}^{\odot}$. For every $p_{j}\in\mathsf{Pr}[\mathbb{P}^{\odot}]$, we define
@@
We now set for every $p_{j}$, $v_{\mathbf{v}}(p_{j})=x$ for some $x\in{\mathcal{V}}_{p_{j}}$. It is easy to see that $v_{\mathbf{v}}(\kappa^{\odot}_{i})=1$ for all $\kappa^{\odot}_{i}\!\in\!\Gamma$ and $v_{\mathbf{v}}(\sigma)\!=\!1$ but $v_{\mathbf{v}}(\tau)\!=\!0$. Thus, $\Gamma,\sigma\not\models_{{\textsf{\L}}}\tau$ which contradicts the assumption that $\sigma\!\in\!\mathcal{S}(\mathbb{P}^{\odot})$. Thus, we have that $\Gamma_{\mathcal{H}},\sigma_{\mathcal{H}}\!\models_{\mathsf{CPL}}\!\tau_{\mathcal{H}}$. It follows now that $\sigma_{\mathcal{H}}$ is a solution to $\mathbb{P}_{\mathcal{H}}$.
@@
For the hardness, we construct a logspace reduction of the solution existence for classical negative Horn abduction problems to the solution existence for Łukasiewicz abduction problems with cover-free theories. Namely, let $\mathbb{P}=\langle\Gamma,\tau,{\mathsf{H}}\rangle$ be a classical Horn abduction problem with $\Gamma$ containing only negative literals and ${\mathsf{H}}\cap\mathsf{Pr}(\tau)=\varnothing$. We define $\mathbb{P}^{\mathsf{C}}=\langle\Gamma^{\mathsf{C}},\tau^{\mathsf{C}},{\mathsf{H}}^{\mathsf{C}}\rangle$ as follows:
$\Gamma^{\mathsf{C}}$ =$=\left\{\bigoplus\limits^{m}_{i=1}(p_{i}\!\leq\!\mathbf{0})\mid\bigvee\limits^{m}_{i=1}\neg p_{i}\in\Gamma\right\}$
$\tau^{\mathsf{C}}$ $=\bigodot_{q\in\tau}(q\!\geq\!\mathbf{1})\odot\bigodot_{\neg q\in\tau}(q\!\leq\!\mathbf{0})$
${\mathsf{H}}$ =$=\{r\!\geq\!\mathbf{1}\mid r\in{\mathsf{H}}\}\cup\{s\leq\mathbf{0}\mid\neg s\in{\mathsf{H}}\}$
It is clear that $\Gamma^{\mathsf{C}}$ is cover-free and that the size of $\mathbb{P}^{\mathsf{C}}$ is linear in the size of $\mathbb{P}$. Now assume that $\sigma\in\mathcal{S}(\mathbb{P})$. We show that $\sigma^{\mathsf{C}}\in\mathcal{S}(\mathbb{P}^{\mathsf{C}})$. First, it is clear that $\Gamma^{\mathsf{C}},\sigma^{\mathsf{C}}\not\models_{{\textsf{\L}}}\bot$ because $\Gamma,\sigma\not\models_{\mathsf{CPL}}\bot$. To check the entailment, observe that there is a unit resolution inference of literals comprising $\tau$ from $\Gamma\cup\{\sigma\}$. It is clear that classical unit resolution is sound for interval clauses: if $v(p\geq\mathbf{1})=1$ and $v((p\!\leq\!\mathbf{0})\oplus\kappa)=1$, then $v(\kappa)=1$. Thus, reusing the classical inference in Ł, we obtain $\Gamma^{\mathsf{C}},\sigma^{\mathsf{C}}\models_{{\textsf{\L}}}\tau^{\mathsf{C}}$, as required. Conversely, let $\sigma\in\mathcal{S}(\mathbb{P}^{\mathcal{H}})$ and define $\sigma_{\mathsf{cl}}=\bigwedge\limits_{(r\!\geq\!\mathbf{1})\in\sigma}\!\!\!\!r\wedge\bigwedge\limits_{(s\!\leq\!\mathbf{0})\in\sigma}\!\!\!\!\neg s$. It is clear that $\Gamma,\sigma_{\mathsf{cl}}\models_{\mathsf{CPL}}\tau$. To see that $\Gamma,\sigma_{\mathsf{cl}}\not\models_{\mathsf{CPL}}\bot$, assume for contradiction that $\Gamma,\sigma_{\mathsf{cl}}\models_{\mathsf{CPL}}\bot$. Then there is a unit resolution inference of the empty clause from $\Gamma\cup\{\sigma_{\mathsf{cl}}\}$. But as we can reuse it in Ł, this would mean that $\Gamma^{\mathsf{C}},\sigma\models_{{\textsf{\L}}}\bot$, contrary to the assumption. Contradiction.
@@
The membership follows from Theorem 15. For the hardness, we reuse the reduction shown in Theorem 15. ∎

### v3

@@ Appendix C Proofs of Section 5
Now let $\mathbf{v}$ be a classical valuation s.t. $\mathbf{v}(\phi)=1$ for all $\phi\in\Gamma_{\mathcal{H}}$, $\mathbf{v}(\sigma_{\mathcal{H}})=1$ but $\mathbf{v}(\tau_{\mathcal{H}})=0$. We need to define an Ł-valuation $v_{\mathbf{v}}$ s.t. $v_{\mathbf{v}}(\lambda)=1$ iff $\mathbf{v}(r_{\lambda})=1$ would hold for every $\lambda$ occurring in $\mathbb{P}$. For every $p_{j}\in\mathsf{Pr}[\mathbb{P}]$, we define
@@
We now set for every $p_{j}$, $v_{\mathbf{v}}(p_{j})=x$ for some $x\in{\mathcal{V}}_{p_{j}}$. It is easy to see that $v_{\mathbf{v}}(\theta)=1$ for all $\theta\!\in\!\Gamma$ and $v_{\mathbf{v}}(\sigma)\!=\!1$ but $v_{\mathbf{v}}(\tau)\!=\!0$. Thus, $\Gamma,\sigma\not\models_{{\textsf{\L }}}\tau$ which contradicts the assumption that $\sigma\!\in\!\mathcal{S}(\mathbb{P})$. Thus, $\Gamma_{\mathcal{H}},\sigma_{\mathcal{H}}\!\models_{\mathsf{CPL}}\!\tau_{\mathcal{H}}$. It follows now that $\sigma_{\mathcal{H}}$ is a solution to $\mathbb{P}_{\mathcal{H}}$.
@@
For the hardness, we construct a logspace reduction of the solution recognition for classical Horn abduction problems to the solution recognition for Łukasiewicz abduction problems with cover-free theories. Namely, let $\mathbb{P}=\langle\Gamma,\tau,{\mathsf{H}}\rangle$ be a classical Horn abduction problem with $\Gamma$ represented as Horn implications $\eta=\bigwedge^{m}_{i=1}p_{i}\rightarrow q_{i}$ and $\theta=\bigwedge^{m^{\prime}}_{i=1}p^{\prime}_{i}\rightarrow\bot$ and ${\mathsf{H}}\cap\mathsf{Pr}(\tau)=\varnothing$. We define
$\eta^{\mathsf{C}}$ $\coloneqq\bigodot\limits^{m}_{i=1}(p_{i}\!\geq\!\mathbf{1})\rightarrow(q\geq\mathbf{1})$ $\theta^{\mathsf{C}}$ $\coloneqq\bigodot\limits^{m^{\prime}}_{i=1}(p^{\prime}_{i}\!\geq\!\mathbf{1})\rightarrow\bot$
and set $\mathbb{P}^{\mathsf{C}}=\langle\Gamma^{\mathsf{C}},\tau^{\mathsf{C}},{\mathsf{H}}^{\mathsf{C}}\rangle$ as follows:
$\Gamma^{\mathsf{C}}$ $=\left\{\eta^{\mathsf{C}}\mid\eta\in\Gamma\right\}\cup\left\{\theta^{\mathsf{C}}\mid\theta\in\Gamma\right\}$
$\tau^{\mathsf{C}}$ $=\bigodot_{q\in\tau}(q\!\geq\!\mathbf{1})\odot\bigodot_{\neg q\in\tau}(q\!\leq\!\mathbf{0})$ (12)
${\mathsf{H}}$ $=\{r\!\geq\!\mathbf{1}\mid r\in{\mathsf{H}}\}\cup\{s\leq\mathbf{0}\mid\neg s\in{\mathsf{H}}\}$
It is clear that $\Gamma^{\mathsf{C}}$ is cover-free and that the size of $\mathbb{P}^{\mathsf{C}}$ is linear in the size of $\mathbb{P}$. Now assume that $\sigma\in\mathcal{S}(\mathbb{P})$. We show that $\sigma^{\mathsf{C}}\in\mathcal{S}(\mathbb{P}^{\mathsf{C}})$. First, it is clear that $\Gamma^{\mathsf{C}},\sigma^{\mathsf{C}}\not\models_{{\textsf{\L }}}\bot$ because $\Gamma,\sigma\not\models_{\mathsf{CPL}}\bot$. To check the entailment, observe that there is a unit resolution inference of literals comprising $\tau$ from $\Gamma\cup\{\sigma\}$. It is clear that classical unit resolution is sound for interval clauses: if $v(p\geq\mathbf{1})=1$ and $v(((p\!\geq\!\mathbf{1})\odot\tau)\rightarrow\lambda)=1$, then $v(\tau\rightarrow\lambda)=1$; if $v(p\leq\mathbf{0})=1$ and $v(\tau\rightarrow(p\geq\mathbf{1}))=1$, then $v(\tau\rightarrow\bot)=1$. Thus, reusing the classical inference in Ł, we obtain $\Gamma^{\mathsf{C}},\sigma^{\mathsf{C}}\models_{{\textsf{\L }}}\tau^{\mathsf{C}}$, as required. Conversely, let $\sigma\in\mathcal{S}(\mathbb{P}^{\mathcal{H}})$ and define $\sigma_{\mathsf{cl}}=\bigwedge\limits_{(r\!\geq\!\mathbf{1})\in\sigma}\!\!\!\!r\wedge\bigwedge\limits_{(s\!\leq\!\mathbf{0})\in\sigma}\!\!\!\!\neg s$. It is clear that $\Gamma,\sigma_{\mathsf{cl}}\models_{\mathsf{CPL}}\tau$. To see that $\Gamma,\sigma_{\mathsf{cl}}\not\models_{\mathsf{CPL}}\bot$, assume for contradiction that $\Gamma,\sigma_{\mathsf{cl}}\!\models_{\mathsf{CPL}}\!\bot$. Then there is a unit resolution inference of the empty clause from $\Gamma\cup\{\sigma_{\mathsf{cl}}\}$. But as we can reuse it in Ł, this would mean that $\Gamma^{\mathsf{C}},\sigma\models_{{\textsf{\L }}}\bot$, contrary to the assumption. Contradiction.
@@
The membership follows from Theorem 15. For the hardness, we reuse the reduction shown in (C). ∎